# Direct comparison of stochastic driven nonlinear dynamical systems for combinatorial optimization

### v2

@@ III.2 Benchmarking results
Oscillator-based Ising machine.— In our experiments we follow Fig. 10 in Ref. Wang and Roychowdhury (2019) and use the parameters used in the aforementioned reference. Tests on small system sizes confirm that the choice of parameters for the OIM is robust for different system sizes. For the OIM the gain is periodic, which helps the system escape local minima by imposing and relaxing the binary constraint repeatedly. While this approach remains untested for other dynamical system, it could potentially improve performance across the different approaches studied in this work. Figure 1 shows that the OIM has the largest TTS values of all method studied for almost all problem sizes studied. We have set the evolution time of the OIM to $20$ steps while the optimal time step is found to be $0.02$, independent of problem size $L$. Thus, the OIM has the same total number of iterations as the Sim-CIM. This indicates that OIM has a lower success rate to find the optimal than, e.g., the Sim-CIM, and we observe larger error bars for $L=18$. It is also possible that the adapted annealing scheme might not be optimal and a more careful tuning of the parameters might improve the TTS.
@@
Simulated bifurcation machine.— For the SBM the gain profile $\alpha$ is a linear function passing through the origin. The other parameters are chosen as follows. $K=\Delta=1$ and $\beta=0.7/\sigma(J)N_{s}$, where $\sigma(\bullet)$ denotes the standard deviation Goto et al. (2019). As pointed out before, the SBM is different from other SDNDSs because it has a second-order equation of motion (Hamiltonian dynamics). Therefore, instead of using the Euler method, which can cause extra numerical errors in solving higher-order ODEs, we apply the symplectic Euler method Goto et al. (2019). Finally, there is no noise term in the SBM, given its origin stemming from adiabatic quantum optimization. In our experiments the parameters of SBM are mostly constant across different system sizes and we only optimize the total number of iterations listed in Tab. 4. From Fig. 1, we observe that the SBM seems to be an effective optimizer for the problems studied, because overall it has one of the smallest TTS for the different system sizes. However, we note that the SBM may require much smaller time steps in some cases.
@@
Hopfield neural network dynamical system.— For the HNN dynamical system we set parameters $\alpha=\beta=1$ and $\phi(x_{l})=\tanh(x_{l}/\zeta(t))$ Caparrós et al. (2002). Specifically, $\zeta(t)>0$ is a geometric function with initial value $\zeta(0)=\zeta_{0}$ and decay rate $\zeta_{d}$. Parameters do not tend to vary across system sizes studied and thus we fix these for all $L$. Remaining parameters are listed in Tab. 4. Note that we simulate a stochastic Hopfield model including a noise term (discrete or continuous), because it has been shown to have better performance than its deterministic counterpart Cai et al. (2020); Liu and Wang (2009). However, in this work we find that such positive effect is only visible for smaller system sizes. Our results show that the HNN dynamical system tends to become trapped in local minima. However, it remains competitive for small to medium system sizes. In terms of the absolute TTS, it is faster than most of the dynamical systems we tested (see Fig. 1) due to its simple dynamics and fast convergence. However, when $L\geq 24$ the HNN dynamical system has difficulties in finding the optimal of the problem studied.
@@
Simulate annealing.— As a benchmark, we also use simulated annealing. In our implementation we use an inverse linear annealing schedule and tune both the decay rate and the total number of iterations to optimal values to determine the TTS. In our experiments we find the final inverse temperature $\beta_{f}=3$ is independent of linear lattice size $L$. The total number of iterations for different system sizes are summarized in Tab. 4. As shown in Fig. 1, the TTS of SA is comparable to the SDNDSs.
@@
Figure 2: Scaling exponent $b$ of different SDNDSs listed in in Table. 1 which corresponds to the slope of the linear fit (in logarithmic scale) to the TTS data shown in Fig. 1. The height of the boxes represents the error bars.
@@ III.3 Scaling exponent
All other dynamical systems, except the OIM discussed below, have comparable scaling exponents (within error bars), which can be ascribed to the similarities between their equations of motion and corresponding dynamics, see Tab. 1. We suspect that the better performance of the Sim-CIM for the problems studied can be ascribed to both larger time steps than other SDNDSs, as well as its unique nonlinear activation, which has also improved the performance of the SBM Goto et al. (2021). However, the improved performance of the Sim-CIM comes at the price of tuning more parameters. Finally, note that the suboptimal performance of the OIM might be due to the fact parameter tuning for this SDNDS is difficult.
@@ IV.1 Effects of numerical integration methods
Algorithms like Sim-CIM, OEO-CIM, and the discrete-time Hopfield neural network are discrete in time. This means that there is no need to specify a numerical integration method. On the other hand, algorithms such as OPO-CIM, GD, and OIM, are first-order ordinary differential equations (ODEs). While the Euler method is the most common choice to integrate these, few efforts have been devoted to investigating how numerical methods affect these optimization routines.
@@
Figure 3(a) shows the scaling results of the Hopfield neural network with different integration methods. The parameters for the Euler method are the same as in Tab. 4. Note that, for RK4, we find that a larger step size ($dt=1$) also works well. For both higher-order integrators (RK4 and RK45) the overall numerical effort is larger than for the Euler method. Given that the scaling exponents agree within error bars, there is no clear indication that using higher-order schemes is needed when integrating the equations of motion.
@@
We further test RK4 for the OIM, because it has relatively complicated dynamics with periodic gains. A comparison between the Euler and RK4 methods is shown in Fig. 4. For $L=4$, $6$, $8$, and $10$ we use $dt=0.05$, whereas for $L\geq 12$ we use $dt=0.02$. Recall that the optimal time step for OIM using the Euler integration scheme is $dt=0.02$, which is independent of system size. Thus, for small problems $L\leq 10$, RK4 can bear a larger time step $dt=0.05$. Again, as can be seen in Fig. 4 the scaling exponent $b$ agrees within error bars and a more sophisticated integrator means only additional numerical overhead.
Summarizing, using different integration schemes does not seem to improve the TTS scaling for both the Hopfield neural network and the OIM. This is in agreement with results for digital memcomputing machines (DMMs)Zhang and Di Ventra (2021) based on memristor circuits. One potential use for RK45 could be for initialization. After a short, burn-in phase an optimal time step is determined, after which the integrator is swapped out by a simple Euler method. Such approach would drastically simplify parameter tuning.
@@ IV.2 Effects of graph connectivity
To test this hypothesis we use the Viana-Bray (VB) model Viana and Bray (1985). For example, for the Viana-Bray model with connectivity 3, each spin is connected to 3 other randomly-chosen spins, i.e., the system has no geometry despite a fixed connectivity. We study a small VB model with $N=2^{7}=128$ variables and successively increase the connectivity in powers of $2$ from $16$ to $128$, where the latter represents the fully-connected Sherrington-Kirkpatrick spin-glass model Sherrington and Kirkpatrick (1975). We use both SA and the SBM to to find optima for this system as a function of connectivity. Data are averaged over 100 disorder realizations of Gaussian disorder with zero mean and standard deviation unity. The results are summarized in Fig. 5. The data clearly show that the TTS increases with connectivity for SA. This is no surprise, because the number of Hamiltonian terms grows with the connectivity and thus the time it takes for one Monte Carlo sweep. In contrast, the TTS decreases for increasing connectivity for the SBM. This suggests that large avalanches are needed for SDNDSs to be effective optimizers. This, in turn, suggests that the denser the graph of the problem to be optimized, the more effective SDNDSs are, and that self-organized critical behavior resulting in large spin rearrangements is critical for optimization in dynamical systems.
@@ V Discussions and conclusions
In this work we have examined several representative combinatorial Ising solvers based on stochastic dynamical nonlinear systems—where binary variables are relaxed to continuous analog spins—and compared them using the same software stack optimizing square-lattice disordered Ising spin-glass problems with planted solutions. A list of the different equations of motion is given in Tab. 1. All dynamical systems share some characteristics, namely a driver term $\alpha$, a global Ising coupling $\beta$, a noise term $g_{l}$, and nonlinear dynamics. In addition, we compared the performance of the driven dynamical systems to simulated annealing as a benchmark. Our results illustrate the advantages and disadvantages some of these stochastic driven nonlinear dynamical systems pose (in comparison to simulated annealing) for a given benchmark problem and moderate system sizes. It would be interesting to do further comparisons for larger system sizes, still leveraging the same software stack. However, this was outside the scope of this project and would require implementations leveraging GPUs. We also analyze the effects of higher-order integration schemes and show that simple Euler dynamics seems, in general, sufficient for most applications.
We also analyzed the effects of graph connectivity on the solver performance. Optimizers require large variable rearrangements to be effective when traversing the search space. Our conjecture that self-organized critical behavior is needed for SDNDS solvers to be effective is backed by numerical results that show that when the connectivity is increased, performance increases despite the number of terms to sum over in the cost function increases as well. In general, most optimization problems are not natively QUBO and also often have constraints that need to be added to the cost function. In the process of mapping optimization problems to QUBOs, as well as including constraints via quadratization into the cost function, the resulting graph tends to be highly- if not fully-connected. This means that optimization heuristics based on stochastic driven nonlinear dynamical systems have an inherent advantage over annealing-based quantum devices Lanting et al. (2014) for real-world applications.
Finally, the dynamical systems listed in Tab. 1 stem from physical phenomena. It would be interesting to investigate if better equations of motion that do not follow physical processes can be found with the goal of improving performance.
@@ Acknowledgements.
H. G. K. would like to thank Austin Hope for inspiration during the final stages of the manuscript.

### v3

@@ III.2.4 Oscillator-based Ising machine
In our experiments we follow Fig. 10 in Ref. Wang and Roychowdhury (2019) and use the parameters used in the aforementioned reference. Tests on small system sizes confirm that the choice of parameters for the OIM is robust for different system sizes. For the OIM the gain is periodic, which helps the system escape local minima by imposing and relaxing the binary constraint repeatedly. While this approach remains untested for other dynamical system, it could potentially improve performance across the different approaches studied in this work. Figure 1 shows that the OIM has the largest TTS values of all method studied for almost all problem sizes studied. We have set the evolution time of the OIM to $20$ steps while the optimal time step is found to be $0.02$, independent of problem size $L$. Thus, the OIM has the same total number of iterations as the Sim-CIM. This indicates that OIM has a lower success rate to find the optimal than, e.g., the Sim-CIM, and we observe larger error bars for $L=18$. It is also possible that the adapted annealing scheme might not be optimal and a more careful tuning of the parameters might improve the TTS.
III.2.5 Simulated bifurcation machine
@@
For the SBM the gain profile $\alpha$ is a linear function passing through the origin. The other parameters are chosen as follows. $K=\Delta=1$ and $\beta=0.7/\sigma(J)N_{s}$, where $\sigma(\bullet)$ denotes the standard deviation Goto et al. (2019). As pointed out before, the SBM is different from other SDNDSs because it has a second-order equation of motion (Hamiltonian dynamics). Therefore, instead of using the Euler method, which can cause extra numerical errors in solving higher-order ODEs, we apply the symplectic Euler method Goto et al. (2019). Finally, there is no noise term in the SBM, given its origin stemming from adiabatic quantum optimization. In our experiments the parameters of SBM are mostly constant across different system sizes and we only optimize the total number of iterations listed in Table 4. From Fig. 1, we observe that the SBM seems to be an effective optimizer for the problems studied, because overall it has one of the smallest TTS for the different system sizes. However, we note that the SBM may require much smaller time steps in some cases.
III.2.6 Hopfield neural network dynamical system
@@
For the HNN dynamical system we set parameters $\alpha=\beta=1$ and $\phi(x_{l})=\tanh(x_{l}/\zeta(t))$ Caparrós et al. (2002). Specifically, $\zeta(t)>0$ is a geometric function with initial value $\zeta(0)=\zeta_{0}$ and decay rate $\zeta_{d}$. Parameters do not tend to vary across system sizes studied and thus we fix these for all $L$. Remaining parameters are listed in Table 4. Note that we simulate a stochastic Hopfield model including a noise term (discrete or continuous), because it has been shown to have better performance than its deterministic counterpart Cai et al. (2020); Liu and Wang (2009). However, in this work we find that such positive effect is only visible for smaller system sizes. Our results show that the HNN dynamical system tends to become trapped in local minima. However, it remains competitive for small to medium system sizes. In terms of the absolute TTS, it is faster than most of the dynamical systems we tested (see Fig. 1) due to its simple dynamics and fast convergence. However, when $L\geq 24$ the HNN dynamical system has difficulties in finding the optimal of the problem studied.
III.2.7 Simulated annealing
@@
As a benchmark, we also use simulated annealing. In our implementation we use an inverse linear annealing schedule and tune both the decay rate and the total number of iterations to optimal values to determine the TTS. In our experiments we find the final inverse temperature $\beta_{f}=3$ is independent of linear lattice size $L$. The total number of iterations for different system sizes are summarized in Table 4. As shown in Fig. 1, the TTS of SA is comparable to the SDNDSs.
@@
Figure 2: Scaling exponent $b$ of different SDNDSs listed in in Table 1 which corresponds to the slope of the linear fit (in logarithmic scale) to the TTS data shown in Fig. 1. The height of the boxes represents the error bars.
@@ III.3 Scaling exponent
All other dynamical systems, except the OIM discussed below, have comparable scaling exponents (within error bars), which can be ascribed to the similarities between their equations of motion and corresponding dynamics, see Table 1. We suspect that the better performance of the Sim-CIM for the problems studied can be ascribed to both larger time steps than other SDNDSs, as well as its unique nonlinear activation, which has also improved the performance of the SBM Goto et al. (2021). However, the improved performance of the Sim-CIM comes at the price of tuning more parameters. Finally, note that the suboptimal performance of the OIM might be due to the fact parameter tuning for this SDNDS is difficult.
@@ IV.1 Effects of numerical integration methods
Algorithms like Sim-CIM, OEO-CIM, and the discrete-time Hopfield neural network are discrete in time. This means that there is no need to specify a numerical integration method. However, algorithms such as OPO-CIM, GD, and OIM, are first-order ordinary differential equations (ODEs). While the Euler method is the most common choice to integrate these, few efforts have been devoted to investigating how numerical methods affect these optimization routines.
@@
Figure 3(a) shows the scaling results of the Hopfield neural network with different integration methods. The parameters for the Euler method are the same as in Table 4. Note that, for RK4, we find that a larger step size ($dt=1$) also works well. For both higher-order integrators (RK4 and RK45) the overall numerical effort is larger than for the Euler method. Given that the scaling exponents agree within error bars, there is no clear indication that using higher-order schemes is needed when integrating the equations of motion.
@@
We further test RK4 for the OIM, because it has relatively complicated dynamics with periodic gains. A comparison between the Euler and RK4 methods is shown in Fig. 4. For $L=4$, 66, 88, and $10$ we use $dt=0.05$, whereas for $L\geq 12$ we use $dt=0.02$. Recall that the optimal time step for OIM using the Euler integration scheme is $dt=0.02$, which is independent of system size. Thus, for small problems $L\leq 10$, RK4 can bear a larger time step $dt=0.05$. Again, as can be seen in Fig. 4 the scaling exponent $b$ agrees within error bars and a more sophisticated integrator means only additional numerical overhead.
Summarizing, using different integration schemes does not seem to improve the TTS scaling for both the Hopfield neural network and the OIM. This is in agreement with results for digital memcomputing machines Zhang and Di Ventra (2021) based on memristor circuits. One potential use for RK45 could be for initialization. After a short, burn-in phase an optimal time step is determined, after which the integrator is swapped out by a simple Euler method. Such approach would drastically simplify parameter tuning.
@@ IV.2 Effects of graph connectivity
To test this hypothesis we use the Viana-Bray (VB) model Viana and Bray (1985). For example, for the Viana-Bray model with connectivity 3, each spin is connected to three other randomly chosen spins, i.e., the system has no geometry despite a fixed connectivity. We study a small VB model with $N=2^{7}=128$ variables and successively increase the connectivity in powers of $2$ from $16$ to $128$, where the latter represents the fully-connected Sherrington-Kirkpatrick spin-glass model Sherrington and Kirkpatrick (1975). We use both SA and the SBM to to find optima for this system as a function of connectivity. Data are averaged over 100 disorder realizations of Gaussian disorder with zero mean and standard deviation unity. The results are summarized in Fig. 5. The data clearly show that the TTS increases with connectivity for SA. This is no surprise, because the number of Hamiltonian terms grows with the connectivity and thus the time it takes for one Monte Carlo sweep. In contrast, the TTS decreases for increasing connectivity for the SBM. This suggests that large avalanches are needed for SDNDSs to be effective optimizers. This, in turn, suggests that the denser the graph of the problem to be optimized, the more effective SDNDSs are, and that self-organized critical behavior resulting in large spin rearrangements is critical for optimization in dynamical systems.
@@ V Discussions and conclusions
In this work we have examined several representative combinatorial Ising solvers based on stochastic dynamical nonlinear systems—where binary variables are relaxed to continuous analog spins—and compared them using the same software stack optimizing square-lattice disordered Ising spin-glass problems with planted solutions. A list of the different equations of motion is given in Table 1. All dynamical systems share some characteristics, namely a driver term $\alpha$, a global Ising coupling $\beta$, a noise term $g_{l}$, and nonlinear dynamics. In addition, we compared the performance of the driven dynamical systems to simulated annealing as a benchmark. Our results illustrate the advantages and disadvantages some of these stochastic-driven nonlinear dynamical systems pose (in comparison to simulated annealing) for a given benchmark problem and moderate system sizes. It would be interesting to do further comparisons for larger system sizes, still leveraging the same software stack. However, this was outside the scope of this project and would require implementations leveraging GPUs. We also analyze the effects of higher-order integration schemes and show that simple Euler dynamics seems, in general, sufficient for most applications.
We also analyzed the effects of graph connectivity on the solver performance. Optimizers require large variable rearrangements to be effective when traversing the search space. Our conjecture that self-organized critical behavior is needed for SDNDS solvers to be effective is backed by numerical results that show that when the connectivity is increased, performance increases despite the number of terms to sum over in the cost function increases as well. In general, most optimization problems are not natively QUBO and also often have constraints that need to be added to the cost function. In the process of mapping optimization problems to QUBOs, as well as including constraints via quadratization into the cost function, the resulting graph tends to be highly if not fully connected. This means that optimization heuristics based on stochastic-driven nonlinear dynamical systems have an inherent advantage over annealing-based quantum devices Lanting et al. (2014) for real-world applications.
Finally, the dynamical systems listed in Table 1 stem from physical phenomena. It would be interesting to investigate if better equations of motion that do not follow physical processes can be found with the goal of improving performance.
@@ Acknowledgements.
We would like to thank Ken Zick for useful comments.